# Editable Concept Bottleneck Models

### v1

@@ Appendix G More Related Work
Influence Function. The influence function, initially a staple in robust statistics [11, 12], has seen extensive adoption within machine learning since [24] introduced it to the field. Its versatility spans various applications, including detecting mislabeled data, interpreting models, addressing model bias, and facilitating machine unlearning tasks. Notable works in machine unlearning encompass unlearning features and labels [37], minimax unlearning [28], forgetting a subset of image data for training deep neural networks [16, 15], graph unlearning involving nodes, edges, and features. Recent advancements, such as the LiSSA method [1, 26] and kNN-based techniques [17], have been proposed to enhance computational efficiency. Besides, various studies have applied influence functions to interpret models across different domains, including natural language processing [18] and image classification [5], while also addressing biases in classification models [36], word embeddings [6], and finetuned models [9]. Despite numerous studies on influence functions, we are the first to utilize them to construct the editable CBM. Moreover, compared to traditional neural networks, CBMs are more complicated in their influence function. Because we only need to change the predicted output in the traditional influence function. While in CBMs, we should first remove the true concept, then we need to approximate the predicted concept in order to approximate the output. Bridging the gap between the true and predicted concepts poses a significant theoretical challenge in our proof.
Appendix H Limitations
It is important to acknowledge that the ECBM approach is essentially an approximation of the model that would be obtained by retraining with the edited data. However, results indicate that this approximation is effective in real-world applications.
Appendix I Broader Impacts
Concept Bottleneck Models (CBMs) have garnered much attention for their ability to elucidate the prediction process through a human-understandable concept layer. However, most previous studies focused on cases where the data, including concepts, are clean. In many scenarios, we always need to remove/insert some training data or new concepts from trained CBMs due to different reasons, such as data mislabelling, spurious concepts, and concept annotation errors. Thus, the challenge of deriving efficient editable CBMs without retraining from scratch persists, particularly in large-scale applications. To address these challenges, we propose Editable Concept Bottleneck Models (ECBMs). Specifically, ECBMs support three different levels of data removal: concept-label-level, concept-level, and data-level. ECBMs enjoy mathematically rigorous closed-form approximations derived from influence functions that obviate the need for re-training. Experimental results demonstrate the efficiency and effectiveness of our ECBMs, affirming their adaptability within the realm of CBMs. Our ECBM can be an interactive model with doctors in the real world, which is an editable explanation tool.

### v2

@@ Appendix H More Related Work
Influence Function. The influence function, initially a staple in robust statistics Cook (2000); Cook & Weisberg (1980), has seen extensive adoption within machine learning since Koh & Liang (2017) introduced it to the field. Its versatility spans various applications, including detecting mislabeled data, interpreting models, addressing model bias, and facilitating machine unlearning tasks. Notable works in machine unlearning encompass unlearning features and labels Warnecke et al. (2023), minimax unlearning Liu et al. (2024), forgetting a subset of image data for training deep neural networks Golatkar et al. (2020; 2021), graph unlearning involving nodes, edges, and features. Recent advancements, such as the LiSSA method Agarwal et al. (2017); Kwon et al. (2023) and kNN-based techniques Guo et al. (2021), have been proposed to enhance computational efficiency. Besides, various studies have applied influence functions to interpret models across different domains, including natural language processing Han et al. (2020) and image classification Basu et al. (2021), while also addressing biases in classification models Wang et al. (2019), word embeddings Brunet et al. (2019), and finetuned models Chen et al. (2020). Despite numerous studies on influence functions, we are the first to utilize them to construct the editable CBM. Moreover, compared to traditional neural networks, CBMs are more complicated in their influence function. Because we only need to change the predicted output in the traditional influence function. While in CBMs, we should first remove the true concept, then we need to approximate the predicted concept in order to approximate the output. Bridging the gap between the true and predicted concepts poses a significant theoretical challenge in our proof.